A charge Q is distributed over two concentric spheres of radii r and R (>r)such that the surface charge density are equal. Find the potential at the common centre.

Grade:12

1 Answers

Saikat Banerjee
6 years ago
It is the surface potential both due to spheres.i t will be added as potential is a scaler quantity..V=kQ/r+kQ/r.K=1/4π€.I hope my answer was useful.